OpenStudy (ny,ny):

Find the range of the function rule y=5x-2 for each domain. {-1.2, 0, 4}

hero (hero):

The domain is x = {-1.2, 0, 4} Plug each value of x into the function individually, to find each corresponding y.

hero (hero):

For example, for x = 0, we substitute zero into the equation to find its y value like so: y = 5x-2 y = 5(0)-2 y = -2 There you go. Good luck.

OpenStudy (anonymous):

|dw:1329007126192:dw1

OpenStudy (ny,ny):

thanks hero